OpenStudy (rissyroo118):

The perimeter of a rectangle is 68 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 9 : 8. Which of the following would be the best equation to use to solve this problem? 2x + 2x = 68 (9 + 8) = 68x 9x + 8x = 68 2(9x) + 2(8x) = 68

OpenStudy (rock_mit182):

i would say d

OpenStudy (rock_mit182):



OpenStudy (wolf1728):

A) 2width + 2length = 68 B) 9w = 8L OR 9W -8L = 0 Multiplying "A" by 4 A) 8W +8L = 272 B) 9W -8L = 0 17W = 272 W = 16 L = 18 Are those equations typed correctly? 'x' is the only unknown in all the equations